# UniCode: A Framework for Generating High-Quality Competitive Coding Problems

### v1

@@ 4.4 Performance Across Algorithmic Paradigms
Strengths in template-driven tasks, weaknesses in complex reasoning.
@@
The models demonstrate high proficiency in deterministic, template-driven tasks such as fundamentals and data structures. These problems, which often involve standard data structure manipulations, are likely well-represented in training corpora from textbooks and online repositories. Their solutions typically follow predictable patterns that models can easily recognize and reproduce. In contrast, performance drops significantly on problems requiring novel reasoning and multi-step planning, such as graph algorithms and dynamic programming problems. These domains often require customized logical deduction rather than simple pattern matching. This performance gap suggests that current models may rely more on memorization of common solutions than on robust reasoning.
5 Related Work
5.1 Competitive Code Generation
Evaluating the code generation capabilities of large language models is a rapidly evolving field, driven by the proliferation of powerful models (Jaech et al., 2024; Li et al., 2023; Guo et al., 2024; Hui et al., 2024; Zhang et al., 2023; Guo et al., 2025; Li et al., 2022; Shao et al., 2024). Foundational benchmarks like HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and APPS (Hendrycks et al., 2021) established evaluation standards based on functional correctness on standalone programming puzzles. However, their static nature and limited difficulty spectrum make them susceptible to data contamination (Oren et al., 2023; Golchin & Surdeanu, 2023; Riddell et al., 2024; Roberts et al., 2023), where models may have seen solutions during pretraining. To address this, subsequent benchmarks (Li et al., ; Gu et al., ; Zhu et al., 2025; Chambon et al., ) draw from competitive programming to assess more complex reasoning, and some mitigate contamination with recent contest problems (Zheng et al., 2025b; Jain et al., 2024), yet updates remain slow. In contrast, our work introduces a scalable framework that generates a virtually infinite stream of novel, competitive coding problems, enabling continuous and contamination-free evaluation.
5.2 Automated Test Case Generation
Generating comprehensive test cases for accurate assessment of algorithmic correctness is difficult to scale because it relies on manual curation (Chen et al., 2021; Hendrycks et al., 2021; Austin et al., 2021; Li et al., ; Quan et al., 2025). Recent efforts leverage large language models for test generation (Chen et al., 2022; Schäfer et al., 2023; Liu et al., 2023a), with frameworks like CodeContests++ (Wang et al., 2025) and LiveCodeBench (Jain et al., 2024) employing this for competitive programming problems. However, these methods often crucially depend on existing solution code (Schäfer et al., 2023; Tufano et al., 2022; Chen et al., 2022; Liu et al., 2023b), making them unsuitable for novel problems. AutoCodeBench (Chou et al., 2025) attempts to solve this by generating problems from existing code; however, the resulting tasks can lack algorithmic complexity.Our framework overcomes these limitations with a stress-driven synthesis pipeline that generates high-quality test cases for novel problems without needing a reference solution.
5.3 Generative Evaluation
Generative evaluation aims to overcome benchmark stagnation and data contamination (Oren et al., 2023; Roberts et al., 2023) by dynamically creating novel tasks, thus assessing a model’s true generalization capabilities rather than its memorization. This paradigm is explored in open-ended game environments, where MCU (Zheng et al., 2025a) combined atomic tasks into complex new ones, and KUMO (Lin et al., 2025) extended this to practical domains like medical diagnosis. Other works apply generative methods to logical reasoning (Parmar et al., 2024), dynamic evaluation on synthetic data structures (Zhu et al., 2023) or agentic tasks (Shi et al., 2025) to achieve more comprehensive analysis. Our work adapts the principle of generative evaluation to assess complex algorithmic code generation. We create diverse coding problems by fusion existing ones (Pei et al., 2025; Huang et al., 2025; Wu et al., 2021) and generating variations to measure true generalization.
6 Conclusion
In this paper, we introduced UniCode, a new framework designed to generate high-quality competitive programming problems. Our key contributions are a biological evolution-inspired generation strategy and a stress-driven test case synthesis pipeline. By combining these approaches, UniCode generates unique and challenging problems alongside robust, high-coverage test suites. We validated UniCode by evaluating 19 state-of-the-art LLMs, confirming its effectiveness as a challenging and highly discriminative benchmark. This work represents a significant step forward in our ability to evaluate generative models in the demanding field of competitive programming.
A limitation of our study is its reliance on a single powerful closed-source LLM, o4-mini, which may introduce bias by generating problems that align with its own strengths.Future work will investigate employing diverse LLMs as problem generators and developing more cost-effective methods.
Ethics Statement
The authors have carefully considered the ethical implications of this work. This work on automated problem generation is guided by the imperative to contribute to society and human well-being by advancing the fair and robust evaluation of AI systems. We have considered potential societal impacts and strived to uphold the highest standards of scientific integrity, transparency, and fairness throughout the research process. Our methodology is designed to avoid harm by promoting the development of more reliable and generalizable AI models, and we have taken care to respect intellectual property and the creative work inherent in algorithmic problem design. This work aligns with the goal of ensuring that progress in AI is measured ethically and contributes positively to the field.
Reproducibility Statement
To support reproducibility, we will publicly release the UniCode benchmark, including all 492 generated problems, test suites with provenance metadata, and detailed prompts used for problem and test case generation. The experimental setup—including model versions, evaluation metrics, and hyperparameters is clearly documented in the paper. Code for the stress driven test synthesis pipeline and problem generation strategies will be provided as supplementary material. Human evaluation protocols and rating guidelines are described in Appendix. All constraints, input/output formats, and tagging taxonomies are fully specified to enable exact replication. We encourage the community to build upon this dynamically generated benchmark for future evaluations.
Appendix A Supplementary Experimental Results
A.1 Performance and cost
This section analyzes the trade-offs between model performance, inference cost, and output verbosity. Figure 5 presents a Unicode leaderboard, plotting the Pass@1 score against the average cost per problem for various models. This visualization highlights a key finding: a higher Pass@1 score often correlates with increased average cost. For example, models like o4-mini (high) and gpt-5 (medium), which are top performers, also incur a relatively higher cost. This chart provides a clear visual representation of the performance-cost efficiency frontier, helping users select models that balance their budget and performance requirements.
Figure 5: Unicode leaderboard. Pass@1 score vs. average cost per problem across various models.
Table 3 provides a detailed breakdown of these metrics. We report the Overall Pass@1 score, the empirical average inference cost per query (in US dollars), and the average number of output tokens. Models are ranked in descending order of their Overall Pass@1 score to emphasize performance. The table is structured to allow for easy comparison between performance, cost, and output length. We have included two distinct groups: reasoning models (top block), which are specifically designed for complex problem-solving, and general-purpose models (bottom block). The specific versions of the models are as follows: gpt-5-2025-08-07, o4-mini-2025-04-16 high, o4-mini-2025-04-16 medium, gemini-2.5-pro, deepseek-v3.1-thinking, deepseek-r1-0528, o3-mini-2025-01-31, qwen3-235b-a22b, gemini-2.5-flash, grok-3-mini, claude-3.7-sonnet:thinking, deepseek-chat-v3.1, gpt-4.1-mini-2025-04-14, gpt-4.1-2025-04-14, qwen3-coder, claude-sonnet-4-20250514, llama-4-maverick:free, gpt-4o-2024-11-20, qwen-2.5-32b-coder, and llama-3.3-8b-instruct.
Model Overall Avg Cost (\$) Avg Output Tokens
o4-mini (high)* 70.3% 0.0269 6016.8
gpt-5 (medium)* 67.7% 0.0390 3940.9
o4-mini (medium)* 66.1% 0.0205 4565.2
google/gemini-2.5-pro* 61.6% 0.2015 20099.6
deepseek-v3.1 (thinking) 60.5% 0.0276 15990.0
deepseek-r1 55.6% 0.0250 11337.5
o3-mini (medium) 55.1% 0.0230 4782.0
qwen3-235b-a22b 53.5% 0.0343 4046.0
gemini-2.5-flash* 47.7% 0.0090 3249.6
grok-3-mini* 46.4% 0.0035 6136.75
claude-3.7-sonnet* 45.5% 0.1282 16856.8
deepseek-chat-v3.1 49.8% 0.0068 3062.5
gpt-4.1-mini* 42.4% 0.0070 3193.2
gpt-4.1* 36.5% 0.0071 877.6
qwen3-coder 35.4% 0.0145 2577.2
claude-sonnet-4* 32.4% 0.0211 1319.8
llama-4-maverick (free) 26.2% 0.0006 1211.4
gpt-4o* 15.4% 0.0139 904.2
qwen-2.5-32b-coder 13.4% 0.0038 898.6
llama-3.3-8b-instruct 5.5% 0.0002 399.4
Table 3: Model ranking by Overall pass@1 (left) with per-model average inference cost and average output length (right). Reasoning models are listed first; closed-source models are marked with an asterisk. Overall percentages are reproduced from the reference table and used to rank models in descending order.
A.2 Performance across Pass@k Settings
@@
The pass rate rises from 66.1% at $k=1$ to 83.5% at $k=10$, an improvement of over 20 percentage points. This indicates that repeated attempts significantly enhance performance, suggesting that many problems require multiple diverse generations to solve. The performance plateaus after $k=8$, implying diminishing returns beyond this point. Even with multiple attempts, the model does not achieve a near-perfect score, illustrating the benchmark has high upper limit and diagnostic value. These results underscore the importance of sampling numerous candidates for difficult tasks and reflect the complexity and variability inherent in the problems.
To analyze the difficulty and recoverability of different problem types, we classified algorithm tags into three groups based on the improvement in pass rates from pass@1 to pass@3: Significant Improvement ($\geq$20%): This group includes tags like tree algorithms, graph algorithms, and mathematical problems. These problems often have multiple correct solutions or implementation approaches, so generating multiple code candidates significantly increases the pass rate. Moderate Improvement (10%–20%):: This group contains tags such as string algorithms and data structures. These problems typically have a more defined algorithmic template, and multiple samples can partially mitigate implementation errors. Minor Improvement ($\leq$10%): This group includes tags like dynamic programming and greedy algorithms. The limited effect of increasing samples is due to either a high initial pass rate or a reliance on structural reasoning rather than random variations. The disparity in improvement across tags indicates that multi-candidate evaluation is highly effective for problems with diverse solution pathways, but yields diminishing returns for tasks requiring deep structural insight or systematic implementation correctness.
@@
A.3 Align with other code benchmarks
To validate UniCode, we conducted a Pearson correlation analysis with LiveCodeBench and LiveCodeBenchPro, using a consistent set of models. As depicted in Figure˜7, we observed a strong positive correlation ($r>0.9$) with LiveCodeBench. Conversely, a strong negative correlation (r<0.9r<-0.9) is found with LiveCodeBenchPro. This negative correlation arises because LiveCodeBenchPro uses ranking-based scores where smaller values indicate better performance, whereas our metric follows a convention in which larger values correspond to better performance. This demonstrates that UniCode provides a consistent and reliable measure of model performance, aligning with established benchmarks.
@@ A.3 Align with other code benchmarks
Appendix B Trustworthy Evaluation with Erroneous Tasks

### v2

@@ A.2 Performance Across Algorithmic Paradigms
The models demonstrate high proficiency in deterministic, template-driven tasks such as fundamentals and data structures. These problems, which often involve standard data structure manipulations, are likely well-represented in training corpora from textbooks and online repositories. Their solutions typically follow predictable patterns that models can easily recognize and reproduce. In contrast, performance drops on problems requiring novel reasoning and multi-step planning, such as graph algorithms and dynamic programming problems, which often necessitate customized logical deduction. This performance gap aligns with our earlier findings: a notable strength in template-driven tasks but a weakness in complex reasoning.
A.3 Performance across Pass@k Settings
@@
The pass rate rises from 66.1% at $k=1$ to 83.5% at $k=10$, an improvement of over 20 percentage points. This indicates that repeated attempts significantly enhance performance, suggesting that many problems require multiple sampling to solve. The performance plateaus after $k=8$, implying diminishing returns beyond this point. Even with multiple attempts, the model does not achieve a near-perfect score, illustrating the benchmark has high upper limit and diagnostic value. These results underscore the importance of sampling numerous candidates for difficult tasks and reflect the complexity and variability inherent in the problems.
To analyze the intrinsic difficulty and recoverability of different problem categories, we group algorithm tags into three classes based on the marginal gain in pass rate from $\text{pass}@1$ to $\text{pass}@3$:
  • Significant Improvement ($\geq 20\%$). This category includes tree algorithms, graph algorithms, and mathematical problems. These tasks often admit multiple valid solution paths or implementation strategies; consequently, sampling multiple candidates substantially increases the probability of success.
  • Moderate Improvement ($10\%\sim 20\%$). This group comprises tags such as string algorithms and data structures. These problems typically follow established algorithmic templates, where additional sampling mainly helps mitigate localized implementation errors or off-by-one bugs.
  • Minor Improvement ($\leq 10\%$). This category features dynamic programming and greedy algorithms. The limited benefit of increased sampling suggests that these problems either have a high baseline pass rate or rely primarily on rigorous structural reasoning rather than implementation variability.
The disparity in improvement across categories indicates that multi-candidate evaluation is highly effective for problems with diverse solution pathways, but yields diminishing returns for tasks that require deep logical reasoning or systemic correctness.
@@
A.4 Alignment with Contamination-Free Code Benchmarks
To verify the validity and robustness of UniCode, we evaluate its alignment with two widely recognized, contamination-free benchmarks: LiveCodeBench and LiveCodeBenchPro. These benchmarks are specifically designed to mitigate the effects of data leakage, providing a reliable gold standard for performance comparison. To ensure experimental integrity, we utilized a consistent ensemble of representative models across all evaluations and conducted a Pearson correlation analysis to quantify the relationship between UniCode and these established metrics.
@@ A.4 Alignment with Contamination-Free Code Benchmarks
The experimental results, as visualized in Figure 9, reveal a high degree of statistical consistency: we observe a strong positive correlation with LiveCodeBench ($r>0.9$), indicating that UniCode’s evaluation results are highly congruent with the scoring mechanisms of LiveCodeBench, and a strong negative correlation with LiveCodeBenchPro (r<0.9r<-0.9), which is expected given the differing scoring conventions—while UniCode follows a “higher-is-better” metric, LiveCodeBenchPro employs a ranking-based system where lower numerical values indicate superior performance.
The high absolute correlation values ($|r|>0.9$) across both benchmarks provide empirical evidence that UniCode serves as a reliable and high-fidelity proxy for model performance. This alignment confirms that our benchmark effectively captures the underlying coding capabilities of models in a manner.
A.5 Performance Stability via Recursive Augmentation
To further investigate the depth and resilience of Large Language Models’ (LLMs) reasoning capabilities, we conducted a recursive augmentation experiment. This process introduces cumulative structural and logical shifts, moving the test cases further away from the data distributions potentially encountered during pre-training.
Experimental Setup
We randomly sampled a subset of 132 problems from Variant-L1 and re-applied the UniCode framework to generate 68 Variant-L2 problems. The evaluation spans three progressive levels:
  • Seed: Original problems curated from human-centered benchmarks (e.g., Codeforces).
  • Variant-L1: First-generation variants generated based on the Seed problems.
  • Variant-L2: Second-generation variants generated by applying UniCode’s augmentation logic to Variant-L1.
Findings and Analysis
As illustrated in ,a consistent performance collapse is observed across all tested models. While models exhibit varying degrees of resilience at L1, the transition to L2 leads to an additional average decline of 7.9%.
Interestingly, the performance drop from L1 to L2 is generally less severe than the initial collapse from Level 0 to L1. We posit that the transition from Seed to L1 primarily serves to decouple statistical shortcuts and rote memorization of canonical training data. Once these shortcuts are neutralized at L1, the subsequent move to L2 tests the model’s intrinsic reasoning depth rather than further memory exploitation.
Specifically, the results reveal a divergence in model robustness: Models like o4-mini demonstrate a resilient reasoning core, with a marginal decline of only 2.6% from L1 to L2. This suggests its internal logic remains stable despite increased complexity. In sharp contrast, Gemini-1.5-Flash experiences a substantial drop of 19.4%, indicating that its performance is highly sensitive to even minor structural perturbations once canonical patterns are removed.
Conclusion
This recursive evaluation confirms that current static paradigms severely overestimate model intelligence by conflating memorization with reasoning. UniCode’s ability to generate deeper-level variants provides a more realistic and rigorous ”upper bound” for evaluating the genuine logical capacities of LLMs.
Figure 10: Robustness Decay under Recursive Augmentation. Accuracy of various LLMs on the original seed problems (Seed), first-order variants (Variant-L2), and second-order variants (Variant-L2) generated by UniCode.
A.6 Code-tag Distribution
Figure 11: Distribution of tags, sub-tags, and skills in UniCode dataset. Best viewed when zoomed in.
A.7 Example Problems
Seed Problem 1: Path of Tasty Dishes
Knowing the devilish nature of Suraj, Ashi is concerned about the Chef and has asked all of Chef’s town people to help. The distance from Chef’s town to Byteland through the path of tasty dishes is $X$ units. They have the location where the magic dishes are and how many people are required to eat it completely. Anyone who eats a dish would go to a long sleep and won’t be able to continue. They have the information about the tribal clans that live along the path of tasty dishes who can be of real help in this journey.
The journey Chef and his friends can be described as follows: There is a total of $B$ dishes on the path. Each dish is located at distance $x_{i}$ ($x_{i-1}<x_{i}$). To minimize the number of friends Chef has to leave behind, all of them have decided that exactly $y_{i}$ of them will eat the $i^{th}$ dish. Also, there are $C$ tribal chef clans. Clan $i$ is located at distance $p_{i}$ ($p_{i-1}<p_{i}$) with a population of $r_{i}$. If a group of at least $q_{i}$ men approaches them, they will join the forces.
Input Format:
  • The first line contains an integer $T$, the number of test cases.
  • Each test case contains:
    • Line 1: $X$ (distance to Byteland).
    • Line 2: $B$ (number of dishes).
    • Line 3: $B$ pairs of space-separated integers $x_{i},y_{i}$.
    • Line 4: An integer $C$, followed by $C$ space-separated triplets $p_{i},q_{i},r_{i}$.
Output Format:
For each test case, print the minimum size of the group (including Chef) needed to reach Byteland.
Constraints:
  • $1\leq T\leq 10,\quad 1\leq X\leq 10^{9},\quad 1\leq B\leq 10000$
  • Subproblem 1 (25 pts): $C=0$
  • Subproblem 2 (75 pts): $1\leq C\leq 10000$
  • $1\leq x_{i}<X,\quad 1\leq p_{i}<X$ (all positions are distinct)
  • $1\leq y_{i},q_{i},r_{i}\leq 10^{14}$
Seed Problem 2: Baratheon’s Reign
Problem Statement: The Baratheons have been ruling in the Seven Kingdoms for many years. King Joffrey Baratheon commanded to build two monuments. The Baratheons have been ruling for $N$ years. Every year is described by an integer $A_{i}$, the level of prosperity.
You are to pick two historical periods $S_{1},F_{1}]$ and $S_{2},F_{2}]$ with the following rules:
  • No overlap: Two periods shouldn’t have common years.
  • Chronological: The first period must start earlier than the second one.
  • Separation: There must be at least $K$ years between $F_{1}$ and $S_{2}$.
Goal: Maximize the total sum of prosperity levels in the chosen periods.
Input Format:
  • Line 1: $T$ (test cases).
  • Each case:
    • Line 1: $N$ (years) and $K$ (gap).
    • Line 2: $N$ integers $A_{1},A_{2},\dots,A_{N}$.
Output Format:
For each test case, output a single line containing the maximum sum.
Constraints:
  • $1\leq T\leq 5,\quad 2\leq N\leq 10^{5},\quad 0\leq K\leq 10^{5}$
  • $10^{9}\leq A_{i}\leq 10^{9},\quad K+2\leq N$
New Problem: Chef’s Grand Expedition
Problem Statement: Chef must journey in two phases.
Phase 1: Recruitment
  • There are $N$ districts in Chef’s town, labeled $1\dots N$. District $i$ has $A_{i}$ potential volunteers ($A_{i}$ may be negative: a negative value means the district actually shuns the effort).
  • Chef may conduct exactly two recruitment campaigns, each on a contiguous interval of districts $[L,R]$. These two intervals must not overlap, and there must be at least $K$ districts between the end of the first and the start of the second.
  • Chef gathers the sum of $A_{i}$ in each chosen interval. His total recruits $H$ is the sum over both intervals (if that sum is negative, he would of course choose intervals giving non-negative sum).
Phase 2: Expedition
  • The path from Chef’s town to Byteland has $B$ magical “dishes” at strictly increasing distances $x_{i}$. To cross dish $i$, exactly $y_{i}$ members of Chef’s party must stop (and thus be lost to sleep).
  • There are $C$ tribal clans at strictly increasing distances $p_{j}$. Clan $j$ will join Chef’s party and contribute $r_{j}$ people, but only if at the moment Chef arrives at $p_{j}$ his current party size is at least $q_{j}$.
Goal: Compute the minimal $G_{0}$ such that Chef can complete Phase 2 alive.
Input Format:
  • Line 1: $T$ (number of test cases).
  • For each test case:
    • Line 1: $N$ and $K$.
    • Line 2: $N$ space-separated integers $A_{1},A_{2},\dots,A_{N}$.
    • Line 3: $B$ (number of dishes).
    • Next $B$ lines: Two integers $x_{i},y_{i}$ for the $i$-th dish.
    • Next line: $C$ (number of clans).
    • Next $C$ lines: Three integers $p_{j},q_{j},r_{j}$ for the $j$-th clan.
Output Format:
For each test case, print one integer: the minimum $G_{0}$.
Constraints:
  • $1\leq T\leq 5,\quad 1\leq N,B,C\leq 10^{5},\quad 0\leq K<N$
  • $10^{9}\leq A_{i}\leq 10^{9}$
  • $1\leq x_{i}<x_{i+1}\leq 10^{9},\quad 1\leq y_{i}\leq 10^{14}$
  • $1\leq p_{j}<p_{j+1}\leq 10^{9},\quad 1\leq q_{j},r_{j}\leq 10^{14}$
Appendix B Technical Implementation and Reproducibility
B.1 Test Cases Quality and Ablation Study
We evaluate test suites using two metrics: correctness (accepting valid solutions) and coverage (rejecting invalid ones). An ideal test suite optimizes both axes simultaneously. Let $S_{\text{correct}}$ and $S_{\text{incorrect}}$ be sets of correct and incorrect submissions. A test suite $M$ passes a submission $s$ if $s$ succeeds on all test cases $m_{i}\in M$, denoted $\text{pass}(s,M)=1$. We define:
$\text{Corr@N}=\frac{|\{\,s\in S_{\text{correct}}\mid\text{pass}(s,M)=1\,\}|}{|S_{\text{correct}}|},$ (1)
$\text{Cov@N}=\frac{|\{\,s\in S_{\text{incorrect}}\mid\text{pass}(s,M)=0\,\}|}{|S_{\text{incorrect}}|}.$ (2)
Table 3: Quality evaluation of generated test suites. We compare our full pipeline against the rStar-Coder baseline (Liu et al., 2025) and perform an ablation study to quantify the contribution of each component. Stage 1 (Brute-Force) uses brute-force solvers on small random inputs. MajVote (Unfiltered Solvers) applies a majority vote to all generated solutions without pre-filtering. MajVote (Filtered Solvers) uses only solutions that passed the Stage 1 stress test. Our Full Pipeline integrates all stages and input types. “Validated suites” is the percentage of problems for which a suite passed all validation checks.
Method Input types Correctness Coverage
Problems with
validated suites (%)
rStar-Coder MajVote (Unfiltered) randrand 86.9% 80.2% 94.3%
Ours Stage 1 (Brute-Force) randrand 91.9% 81.5% 98.2%
MajVote (Unfiltered) allall 86.7% 85.2% 93.9%
MajVote (Filtered) allall 93.8% 84.3% 92.8%
Full Pipeline allall 94.5% 86.0% 94.8%
Setup.
We evaluated 80 problems from the Test-Eval dataset (Yang et al., 2025b), each having an average of 200 pass/fail solutions, using test suites of size $N=50$ (see Section B.2 for test suite composition). Note that our approach to majority voting differs from rStar-Coder in its granularity. While rStar-Coder aggregates at the solution level and discards a problem unless a majority of solutions exhibit identical input–output behavior; our aggregation is performed per test case, so disagreement on individual cases does not invalidate the entire problem.
Results and Ablation Analysis.
As shown in Table 3, our full pipeline significantly outperforms the baseline in both correctness (94.5% vs. 86.9%) and coverage (86.0% vs. 80.2%). This performance gain is dissected through a series of ablation experiments:
  • Impact of Stage 1 Filtering: Comparing MajVote (Unfiltered) with MajVote (Filtered) reveals that our brute-force validation on small inputs is pivotal for correctness, yielding a +7.1% absolute improvement (86.7% to 93.8%) by eliminating erroneous solver candidates.
  • Contribution of Adversarial Inputs: The transition from Stage 1 (random inputs) to the Full Pipeline highlights the role of adversarial and corner inputs in enhancing coverage (increasing from 81.5% to 86.0%), proving their efficacy in exposing complex-case failures.
  • Robustness of Aggregation: Despite the increased complexity, our per-test-case adjudication maintains a high “validated suite” rate (94.8%), outperforming rStar-Coder’s rigid solution-level majority vote.
While the automatic generation system can not be error-free, our analysis in App. B.5 confirms that the resulting benchmark remains statistically reliable for evaluating code generation.
B.2 Test-Suite Composition and Parameter Selection
In this section, we provide a detailed justification for the composition of the final test suite $S$, which consists of 50 test cases with a fixed distribution: 20 random ($I_{\text{rand}}$), 20 adversarial ($I_{\text{adv}}$), and 10 corner ($I_{\text{corn}}$) inputs. This configuration was determined through an extensive empirical evaluation aimed at balancing correctness and coverage.
Experimental Setup
We conducted a hyperparameter sweep over multiple test suite compositions, evaluating each configuration on a held-out set of 48 problems with 960 human-crafted solutions. Each configuration was assessed using two key metrics:
  • Correctness: the proportion of valid solutions that pass all test cases.
  • Coverage: the proportion of invalid solutions that are correctly rejected.
Table 4: Representative sweep results across different distributions.
Distribution Correctness (%) Coverage (%)
(5, 5, 0) 97.9 77.0
(10, 10, 5) 95.8 81.4
(20, 20, 10) 94.0 87.5
(30, 30, 20) 91.7 88.7
(50, 50, 20) 91.7 90.0
We observe that smaller test suites (5,5,0) achieve high correctness but suffer from low coverage, failing to detect many faulty solutions. Larger suite such as (30,30,20), improves coverage marginally but at the cost of increased sensitivity to corner test cases and higher computational cost. The configuration (20, 20, 10) is selected as the optimal ”elbow point. It maintains high correctness 94.0% while achieving broad coverage 87.5%, providing a rigorous filtering mechanism without being prohibitively expensive or overly punitive to valid code.
B.3 Analysis of Generator Bias
A potential concern in generative evaluation is the risk of “generator bias”, where a model performs better on problems it generated itself. To address this concern, we performed validation using an alternative, open-source generator deepseek-r1 to generate a new set of 104 problems across 5 distinct tags. We then benchmark 6 models of varying capability levels on this independently generated set and compared the results to their performance on the standard o4-mini-generated UniCode problems and human-curated no data contamination LiveCodeBench [^2] Initial Release: 8/1/2024 to 5/1/2025.
Model Unicode (deepseek-generated) Unicode (o4mini-generated) LiveCodeBench (human-curated, no contamination)
gpt-5 72.5% 67.7%
o4-mini 70.2% 66.9% 74.2%
deepseek-r1 61.6% 56.6% 73.1%
o3-mini 51.0% 57.1% 63.0%
gpt-4.1-mini 41.3% 42.4% 53.2%
gemma-3-27b-it 14.6% 13.0%
Table 5: Pass@1 rates across different problem generators. While absolute scores show fluctuation across the three datasets, indicating differences in generator difficulty or stylistic preferences. The relative model hierarchy remains highly consistent (Pearson $r=0.984$ between deepseek-r1 and o4-mini generated sets).
As indicated in Table 5, the results do not support the presence of significant self-preference bias. The relative ranking of models remains highly consistent across all three datasets, with a Pearson correlation of $r=0.984$ ($p=4\times 10^{-4}$) between model performances on the deepseek-r1-generated and o4-mini-generated problem sets.
Although deepseek-r1 performs markedly better on its own problems than on o4-mini’s (61.6% vs. 55.6%), a similar performance gap persists on human-curated LiveCodeBench. This suggests the difference stems from stylistic preferences inherent to each generator, rather than intentional bias. Therefore, to mitigate any residual stylistic bias and ensure robust evaluation, we will release UniCode-Multi benchmark, a composite benchmark aggregating $100$ problems each from o4-mini, gpt-5, gemini-2.5-pro, Deepseek-r1, and qwen3-235b-a22b models.
B.4 Validity of Generated Problems
To rigorously assess the utility and complexity of our generated benchmark, we conducted an extensive human evaluation involving 113 problems (approx. 23% of the total problem set). This process required over 30 hours of expert labor, as each problem underwent 10–20 minutes of in-depth analysis by competitive programming veterans.
Expert Annotation Protocol
We recruited five independent annotators: senior competitive programmers and algorithm engineers with over 5 years of experience and a Codeforces rating of 2100+ (Master level or above). To ensure objectivity, we employed a blinded rating protocol to evaluate the problems:
  • Validity & Rigor. The evaluation yielded a high Validity rate of 98.2%. Technical analysis of the few invalid cases (1.8%) revealed that they were primarily due to minor output specification ambiguities rather than fundamental logical flaws or insurmountable constraints.
  • Inter-Annotator Agreement. We observed a 92.3% agreement rate among the five experts. This level of consensus, especially given the problems’ complexity, ensures the clarity and formal precision of the generated statements.
Figure 12: Human rating website. Best viewed when zoomed in.
Figure 13: Human rating website. Best viewed when zoomed in.
B.5 Trustworthy Evaluation with Erroneous Tasks